OpenStudy (anonymous):

A water tank 3m in diameter and 6m high is made from a steel taving a thickness of 12mm. When tank is filled w/water determine the circumference stress. Det. the tank stressL at the buttom of the tank when it is filled w/water. if the circumference stress is limited to 5Mpa what is the max. height of the water?

OpenStudy (jack1):

Volume of tank = ? @patbigFat

OpenStudy (anonymous):

not stated sir ..

OpenStudy (jack1):

nono... can u work it out please?

OpenStudy (jack1):



OpenStudy (anonymous):

V=Pi x r ^2 x h ?

OpenStudy (jack1):

yup

OpenStudy (anonymous):

is it 169.6464m?

OpenStudy (jack1):

no

OpenStudy (jack1):

what are u using for r?

OpenStudy (anonymous):

raduis sir?

OpenStudy (jack1):

yeah... what number are u plugging into ur equation tho... for r?

OpenStudy (anonymous):

3 sir?

OpenStudy (jack1):

no... 3 = diameter radius is half of diameter

OpenStudy (jack1):

so give it another shot with r = 1.5 so V=??

OpenStudy (anonymous):

56.54 m3

OpenStudy (jack1):

i got 42.41... how are u calculating this? 3.14159 x 1.5^2 x 6 =42.4115

OpenStudy (anonymous):

my mistake sorry ..

OpenStudy (jack1):

all good dude now... surface area of inside of tank? = ??

OpenStudy (anonymous):

i tot when i sqr 1.5 = 3 so i did a manual calculating .... my bad....

OpenStudy (anonymous):

SA = 2 × pi × r2 + 2 × pi × r × h ?

OpenStudy (anonymous):

70.686 m^2 ?

OpenStudy (anonymous):

guyz help?

OpenStudy (irishboy123):

hoop stress \(\large \sigma_{\theta} = \frac{P \ D}{2t}\) D = 3, t = 12mm get pressure from bernoulli \(P = \rho g h\) h = 6 plug and play [if you know some calculus, the formula is deriveable; but i think it is also pretty standard for thin walled pressure vessels. been ages since i did this stuff.] you can get steel's material properties online, for the end bit you will need its tensile strength and the hoop stress is measured around the circumference , ie tangentially to the circle

OpenStudy (irishboy123):

actually you don't need to look at steel at all you are told "the circumference stress is limited to 5Mpa" so you are now solving h for a given \(\sigma_{\theta}\)

OpenStudy (anonymous):

Uhmm.m P=pgh by the way what's "pg" ?

OpenStudy (irishboy123):

\(\rho \equiv\) density of fluid, here water g \(\equiv\) gravity = \(9.8 m/s^2\)

OpenStudy (jack1):

sorry guys, fell asleep last night

OpenStudy (jack1):

howd u go with this?